Repairs Common to Both Units

1997 Aug: Power supply, audio, and metering boards reverse-engineered to produce schematic.
1997 Aug: Repairs:
  • CPU board and housing mountings were very awkward. Redone.
  • Display board was mounted on flimsy plastic posts and bracket from back corner to heat sink. Replaced with single solid bracket.
  • Door hinges reglued.
  • Original heat sink ran very hot. Replaced with larger separate units.
  • Various connector pins had loosened and become intermittent. Resoldered.
  • Display board +5/GND wiring redone.
  • To avoid replacing battery with NiCd, memory power supply modified for use with alkaline batteries.
2012 Jun: Reverse engineering to produce schematic completed.
2012 Jun: 74S471 and 1831 ROMs dumped from both units and compared. Match.
2012 Jun: Fixed-soldered twisted-pair wiring between CPU board and RF board replaced with wiring incorporating connectors so CPU board may be removed without desoldering.
2012 Jun: Wiring between rotator optical sensors and CPU board replaced with slightly longer wires so plug can be removed with CPU board in place.


Unit 71374

Year of Manufacture: 1977 based on component dates codes.
Date of Receipt: 1997 Aug (for JP)
State upon Receipt: Poor condition due to corrosion from leaking NiCd batteries.
1997 Aug: Repairs:
  • Lamp board brackets cut down.
  • Lamp board trimmed to fit better.
  • Lamp board connector wiring replaced.
  • IF board raised at one end with washers to keep contacts from shorting to chassis.
  • Old style of multiplex board did not have provision for mute signal being fed from the metering board to the PS board. The manufacturer dealt with this by soldering wires between the boards. Replaced this with a new connector with an additional pin between ME3 and SA1, and a pin to plug into SA4.
  • Secondary mute circuitry was tacked on between the multiplex board and a terminal strip on the heat sink. Removed, outputs rewired, new cable to output jacks installed.
  • Stereo lamp replaced.
2012 Jun: Selectivity switch loose due to fractured solder joints. Resoldered.
2013 Jan: Front panel backlight incandescent lamps are all burned out. Uses eight 6V tubular bulbs, or could use 16-17 grain-of-wheat bulbs.
2013 Jan: Front panel removed for light polishing with Silvo to remove what appears to be slight spot corrosion on aluminum.
2013 Jan: Erratic activity from microprocessor system: even channels sometimes show up while using tuning knob, call sign disappears after being entered, tuning bumps a channel or two sporadically. Pin 40 of 1802 should be connected to +5 but is O.C. Bad socket pin (cheap molex pin socket). Plastic top of socket pried off, pins retensioned. Connection restored, unit functioning. Should replace socket.
2013 Aug: Loss of signal after a few minutes of operation. Tuning/display still working. 680 ohm dropping resistor in 30 V supply burned up. 10@35V tantalum cap shorted. Cap replaced with 15@50V, resistor replaced.
2013 Aug: 3*AAA battery holder installed. Memory line supply disconnected and diode rerouted to V+5.
2013 Aug: V+15 heat sink replaced with better type. Noted that regulator receives 27V at input, and that STEREO lamp at 100mA results in significant wasted power in regulator.
2016 Mar: When rotating dial the up the frequency sometimes goes a channel down, and sometimes switches to even channels. No obvious problem source.
2016 Apr: Tuning direction error of Mar seems to be gone. Dial lamps still need doing. Unit returned to Jurgen.


Unit 71507

Year of Manufacture: 1977 based on component dates codes.
Date of Receipt: 1997 Aug
State upon Receipt: Poor condition due to corrosion from leaking NiCd batteries. Power supply board mostly destroyed. Portion of CPU board shows corrosion.
1997 Aug: Repairs:
  • Chassis stripped and wirebrushed to remove corrosion from leaking NiCd battery.
  • Assorted corroded hardware wirebrushed or replaced.
  • Left door hinge screw hole tapped and new screw installed.
  • Touch sensor board trimmed to fit better.
  • Friction spring added to tuning knob.
  • 100F cap on IF board had loosened. Resoldered.
  • 1N457 diode on IF board was tacked on to solder side of board. Holes drilled to move diode to component side.
  • CPU to Front End wiring replaced.
  • Power supply and associated wiring seriously corroded from leaking NiCd battery. Rebuilt.
  • Power switch would not latch. Fixed.
  • Touch sensor pins had been butt-soldered to the board and loosened as a result. Pins moved in to give greater solder support and resoldered.
  • "Narrow Band Adjust" 10K trimmer on metering board had a manufacturing flaw which resulted in a partial open to one leg and hence the outputs were permanently muted. Replaced.
  • Some LEDs in both the numeric and linear display had loosened and connections become intermittent. Resoldered.
  • "LEFT" touch sensor responded intermittently. Associated 555 IC replaced.
2012 Jun: One LM703 in IF stages observed to have an 8.5pF ceramic capacitor tacked on between the input and output (UIF3, pins 1 & 5).
2012 Jun: Setting alpha call-sign does not function properly. Once in alpha set mode, rotating the selector knob to select a symbol changes both the symbol and the position. Action does not seem entirely deterministic.

Address lines from UCP11 to UCP10 not going low with A0/A1 when corresponding p2K/p1K is high. Traced to faulty UCP11 not deselecting A inputs quickly enough. UCP11 replaced. Solved.

2012 Jul: Tuning knob rebrushed to remove corrosion.
2012 Jul: Heat sink on rightmost 9368 on display board loose, reglued with crazy glue. Doesn't heat up as much as others, glue not thermally conductive?
2012 Jul: L/R variable outputs observed to have been wired such that the outputs are swapped. Labels on PCB (new version of audio board) are in error. Outputs rewired to correct.
2012 Jul: Resistors at output of multiplex filter cans observed to be different (L=6.8K,R=6.2K). 6.2K replaced so both are 6.8K.
2012 Jul: Bottom shield for CPU board not reinstalled, screwed to chassis bottom plate. Will need trimming for depth, or CPU board must be raised further, if shield is to be reinstalled.
2013 Jan: Front panel removed for light polishing with Silvo to remove what appears to be slight spot corrosion on aluminum.
2013 Aug: Meters are dead, loss of signal in narrow-band IF mode, audio still present in wideband mode. On UF1 307 IC, pin 7 has 0V. Internal 100-ohm resistor between pins 6 and 7 measures open. Replaced with 110-ohm (other 307s measure 120-130).
2013 Aug: Memory line supply disconnected and diode rerouted to V+5.
2013 Aug: Stereo lamp not functioning. 2N3702 drive transistor BE open. Note pinout is ECB not EBC. Replaced with 2SA573. Gain is a little low but affecting lamp current by < 1mA.



  Log
Micro/CPU-100
bhilpert
2012 Jun